OpenStudy (anonymous):

How do I solve: Find two numbers whose sum is 25 and the sum of whose reciprocals is 1/6. Can anyone help me with this problem? So far I have, x + y = 25 and 1/x + 1/y =1/6

OpenStudy (anonymous):

what do you mean by thro ?

OpenStudy (anonymous):

x+y=25, so x = 25 -y 1/x + 1/y = 1/6 From the second equation, you get (y+x)/xy = 1/6 6x + 6y = xy Substitute x into the equation, 6(25 -y) + 6y = y(25 -y) 150 - 6y + 6y = 25 y - y^2 y^2 - 25 y + 150 = 0 (y -10) (y-15) = 0 therefore y =10 or 15, and x = 15 or 10 (sorry about before)

OpenStudy (anonymous):

No problem, thank you so much! I'm awful at math.